Calculate the Volume occupied by 5.25g of Nitrogen at 26 degree cecius and 74.2cm of pressure

Grade:9

2 Answers

Kartikey sharma
10 years ago

P= 74.2/76 atm

V=?

m=5.25 g

M = 28 g

T=273+26

T=299 K

PV=nRT

74.2/76 x V= m/M x  0.0821 x 299

V=5.25 x 0.0821 x 299 x 76 /74.2

V=4.71074201 L

 

Ayush kumar
5 years ago
We know that PV=nRT
So V=nRT
            P
For n
n= w/M
So 5.35/28.    (Mol. Wt. Of N2 is 28) and when it is not given then we take their gas not atom
T is taken in Kelvin. So 273+26=299
And pressure in atm 7.42/76
V=5.35×0.082×299×76.  
            74.2
=4.71.Answer
 
 
 
 
 
 
 
Answer=4.71 and this question is from book written by RC mukherjee pg no 16 qno 11